Solve the system by any method. WRITE ANSWER(S) AS ORDERED PAIR(S). 7) y = x2 - 3 y = -2x - 3
Answer #1

solution - 7. Y = x 3 - 0 Y = -2x-3 - from Equation & Equation ② x-3 = -22-3 x+2x20 x(x+2) =0 xzol x+2=0 x=-2 put x=0 then Y=